# Krylov Complexity in early universe

### v1

@@ Abstract
The Lanczos algorithm offers a method for constructing wave functions for both closed and open systems based on their Hamiltonians. Given that the entire early universe is fundamentally an open system, we apply the Lanczos algorithm to investigate Krylov complexity across different phases of the early universe, including inflation, the radiation-dominated period (RD), and the matter-dominated period (MD). Notably, we find that Krylov complexity differs between the closed and open system approaches. To effectively capture the impact of potentials during the RD and MD phases, we analyze various inflationary potentials—including the Higgs potential, the $R^{2}$ inflationary potential, and chaotic inflationary potential—taking into account the violations of slow-roll conditions. This analysis is conducted in terms of conformal time through the preheating process. Our numerical results indicate that the evolution of Krylov complexity and Krylov entropy is remarkably similar across both methods, regardless of the potential under consideration. Additionally, we rigorously construct what is referred to as an open two-mode squeezed state, utilizing the second kind of Meixner polynomials. Based on this construction, we are the first to calculate the evolution equations for $r_{k}$ and $\phi_{k}$ as they relate to the scale factor. Our findings suggest that dissipative effects lead to a rapid decoherence-like behavior. Moreover, our results indicate that inflation behaves as a strongly dissipative system, while both the radiation-dominated and matter-dominated phases exhibit characteristics of weak dissipation. This research provides new insights into exploring the universe from the perspective of quantum information.
arxiv: 1234.56789
@@ 1 Introduction
Recently, the Krylov complexity has been broadly applied to condensed matter and high-energy physics. Ref. Muck:2022xfc has systematically analyzed the Krylov complexity in various polynomials dubbed as the orthogonal basis. Due to the importance of the SKY model in high energy physics, the Krylov complexity can also be implemented into this field Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk . Meanwhile, the Krylov complexity can be applied to the generalized coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye and even extend into the quantum field theory He:2024hkw ; He:2024xjp . Ref. Dymarsky:2019elm proved that the chaos can be understood as a delocalization in Krylov complexity. Furthermore, one could observe the exponential growth of Krylov complexity n integrable systems with saddle-dominated scrambling, whose universal features can be verified by Huh:2023jxt . More recent developments of Krylov complexity can be found in Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj . Especailly, the concept of Krylov complexity can be extended into in various fields Nandy:2024htc .
@@ 3.2 Various inflationary potentials
Fig. 2 depicts the evolution of $V_{,\phi\phi}$ with respect to $\log_{10}a$ in RD and MD, using the third iteration of $V_{2}$ and $V_{3}$ via Fig. 1.The evolution of inflation is neglected due to the violation of slow-roll conditions. From Fig. 2, we could see that the effective mass of $R^{2}$ inflation will decrease in RD and then it will grow to the constant ($10^{-12}$) with the oscillations. For the potential of the Higgs field, it is different which shows that there is a peak in RD and after that, it will decrease into some constants. But, it indicates that the effective mass of $V_{2}$ and $V_{2}$ are of almost the same order based on the COBE normalization Planck:2015sxf .
@@ 4.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
Figure 3: The numerical solution of $R_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.. The effect of potential is always ignored in slow-roll inflation. But in RD and MD periods, the influence can be seen by using different potential energies as shown in these panels.
Before analyzing the figure of $r_{k}$, it’s important to note that the Krylov complexity depends only on $r_{k}$. Therefore, we only provide the numerical values of $r_{k}$ in Fig. 3 based on Eq. (37). In this equation, we have set $H_{0}=1$ for simplicity and used various values of $k$ in different time periods. The first panel of Fig. 3 illustrates the evolution of $r_{k}$ during inflation. There is an oscillation before the horizon exits, and then $r_{k}$ will grow linearly, consistent with the standard case of Li:2024kfm . As we transition to the radiation-dominated era (RD), we observe that $r_{k}$ will increase to some constant value that oscillates around this value. Interestingly, different potentials produce almost the same trend for $r_{k}$, indicating that the specific potential does not significantly impact the evolution of $r_{k}$. This observation also applies to the matter-dominated era (MD), where the evolution trend is similar to RD. However, we notice that varying momentum significantly affects $r_{k}$, leading to oscillations as the value of $k$ decreases. Although we cannot distinguish between the various potentials using Fig. 3, our numerical data still demonstrates that the evolution of $r_{k}$ is significantly different whether there is a potential or not. It’s worth noting that we cannot provide the values of $r_{k}$ without a potential, as this is not physically meaningful. The potential cannot be ignored due to the violation of slow-roll conditions.
In this section, we have analyzed the evolution of $r_{k}$ for different inflationary potentials within a single-field inflationary framework. Our research indicates that the value of $r_{k}$ increases during inflation. The behavior of $r_{k}$ in radiation-dominated (RD) and matter-dominated (MD) eras is similar and will oscillate around constant values. Our numerical analysis also shows that various inflationary models produce a similar trend for $r_{k}$. Additionally, we have observed that the scale factor and momentum have a significant impact on the evolution of $r_{k}$. All of the details are encoded in Fig. 3.
@@ 5.1 Krylov complexity
Figure 4: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
@@ 5.2 Krylov Entropy
Figure 5: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
@@ 6.1.1 $b_{n}$, $c_{n}$ and $\mu_{2}$
Figure 6: The plot of $b_{n}=\lambda$ with $n=2$.The numerical of $b_{n}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
@@
Figure 7: The numerical of dissipative coefficient $\mu_{2}$ in terms of $\log_{10}a$ for RD and MD, where we set $H_{0}=1$, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity. The value of $mu_{2}$ indicates the RD and MD are weak dissipative systems.
@@ 6.2 General discussion of $\phi_{n}$
where we have defined $\tilde{c}_{n}=n\chi\mu$ and $b_{n}=n\alpha$. As $\eta$ approaches infinity, the solution of Eq. (52) has the form $\phi(n)\propto e^{-n\xi}$ with $\xi=\frac{2\alpha}{\mu\chi}$, where this form will appear in RD or MD due to their weak dissipative nature. In the weak dissipative region, the corresponding Krylov complexity behaves as follows: it increases exponentially at the onset, with the average position given by $C_{k}\propto e^{2\alpha\eta}$ before reaching $\eta_{*}$. After this point, the Krylov complexity saturates to a constant value $\xi$ when $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln{\frac{2\alpha}{\chi\mu}}$. It is straightforward to demonstrate that $\exp(2a\eta_{*})=\xi$. After some simple algebra, we could obtain the formula $\xi$ in our case as follows,
@@ 6.3 Wave function within the method of open system
$\sum_{n=0}^{\infty}M_{n}(x;\delta,\eta_{1})\frac{t^{n}}{n!}=((1+\delta t^{2})+t^{2})^{-\eta/2}\exp(x\arctan(\frac{t}{1+\delta t}))$ (58)
@@
$|\mathcal{O}(t))=e^{i\mathcal{L}\eta}|e_{0})=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\sum_{n=0}^{\infty}|1-\mu_{2}^{2}|^{\frac{n}{2}}\frac{(-1\exp(2i\phi_{k}(\eta))\tanh\eta)^{n}}{(1+\mu_{2}\tanh\eta)^{n}}|e_{n})$
@@ 6.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open system’s method
Figure 8: The numerical of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
@@ 6.6 Krylov complexity and Krylov entropy with the approach of open system
Figure 9: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=0.01$ at period inflation and RD, $k=0.005$ at period MD for simplicity.
Figure 10: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=0.01$ at period inflation and RD, $k=0.005$ at period MD for simplicity.

### v2

@@ Abstract
The Lanczos algorithm offers a method for constructing wave functions for both closed and open systems based on their Hamiltonians. Given that the entire early universe is fundamentally an open system, we apply the Lanczos algorithm to investigate Krylov complexity across different phases of the early universe, including inflation, the radiation-dominated period (RD), and the matter-dominated period (MD). Notably, we find that Krylov complexity differs between the closed and open system approaches. To effectively capture the impact of potentials during the RD and MD phases, we analyze various inflationary potentials, including the Higgs potential, the $R^{2}$ inflationary potential, and chaotic inflationary potential, which is taking into account the violations of slow-roll conditions. This analysis is conducted in terms of conformal time through the preheating process. Our numerical results indicate that the evolution of Krylov complexity and Krylov entropy is remarkably similar within distinctive potentials in RD and MD. Additionally, we rigorously construct what is referred to as an open two-mode squeezed state, utilizing the second kind of Meixner polynomials. Based on this construction, we are the first to calculate the evolution equations for $r_{k}$ and $\phi_{k}$ as they relate to the scale factor. Our findings suggest that dissipative effects lead to a rapid decoherence-like behavior. Moreover, our results indicate that inflation behaves as a strongly dissipative system, while both the radiation-dominated and matter-dominated phases exhibit characteristics of weak dissipation. This research provides new insights into exploring the universe from the perspective of quantum information.
@@ 1 Introduction
Recently, the Krylov complexity has been broadly applied to condensed matter and high-energy physics. Ref. Muck:2022xfc has systematically analyzed the Krylov complexity in various polynomials dubbed as the orthogonal basis. Due to the importance of the SKY model in high energy physics, the Krylov complexity can also be implemented into this field Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk . Meanwhile, the Krylov complexity can be applied to the generalized coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori ; Kundu:2023hbk , topological phases of matter Caputa:2022eye and even extend into the quantum field theory He:2024hkw ; He:2024xjp . Ref. Dymarsky:2019elm proved that the chaos can be understood as a delocalization in Krylov complexity. Furthermore, one could observe the exponential growth of Krylov complexity in integrable systems with saddle-dominated scrambling, whose universal features can be verified by Huh:2023jxt . More recent developments of Krylov complexity can be found in Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj ; Balasubramanian:2024ghv ; Bhattacharya:2024hto ; Mohan:2023btr . Especailly, the concept of Krylov complexity can be extended into in various fields Nandy:2024htc .
@@ 3.2 Various inflationary potentials
Fig. 2 depicts the evolution of $V_{,\phi\phi}$ with respect to $\log_{10}a$ in RD and MD, using the third iteration of $V_{2}$ and $V_{3}$ via Fig. 1.The evolution of inflation is neglected due to the violation of slow-roll conditions. From Fig. 2, we could see that the effective mass of $R^{2}$ inflation will decrease in RD and then it will grow to the constant ($10^{-12}$) with the oscillations. For the potential of the Higgs field, it is different which shows that there is a peak in RD and after that, it will decrease into some constants. But, it indicates that the effective mass of $V_{2}$ and $V_{3}$ are of almost the same order based on the COBE normalization Planck:2015sxf .
@@ 4.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
Figure 3: The numerical solution of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD, $k=0.005$ at MD for simplicity. The effect of potential is always ignored in slow-roll inflation. But in RD and MD periods, the impact of potential is taken into account.
Before analyzing the figure of $r_{k}$, it’s important to note that the Krylov complexity depends only on $r_{k}$. Therefore, we only provide the numerical values of $r_{k}$ in Fig. 3 based on Eq. (37). In this equation, we have set $H_{0}=1$ for simplicity and used various values of $k$ in different time periods. The first panel of Fig. 3 illustrates the evolution of $r_{k}$ during inflation. There is an oscillation before the horizon exits, and then $r_{k}$ will grow linearly, consistent with the standard case of Li:2024kfm . As we transit to RD, we observe that $r_{k}$ will increase to some constant value that oscillates around this value. Interestingly, different potentials produce almost the same trend for $r_{k}$, indicating that the specific potential does not significantly impact the evolution of $r_{k}$. This observation also applies to MD, where the evolution trend is similar to RD. However, we notice that varying momentum significantly affects $r_{k}$, leading to oscillations as the value of $k$ decreases. Although we cannot distinguish between the various potentials using Fig. 3, our numerical data still demonstrates that the evolution of $r_{k}$ is significantly different whether there is a potential or not. It’s worth noting that we cannot provide the values of $r_{k}$ without a potential, as this is not physically meaningful. The potential cannot be ignored due to the violation of slow-roll conditions.
In this section, we have analyzed the evolution of $r_{k}$ for different inflationary potentials within a single-field inflationary framework. Our research indicates that the value of $r_{k}$ increases during inflation. The behavior of $r_{k}$ in RD and MD eras is similar and will oscillate around constant values. Our numerical analysis also shows that various inflationary models produce a similar trend for $r_{k}$. Additionally, we have observed that the scale factor and momentum have a significant impact on the evolution of $r_{k}$. All of the details are encoded in Fig. 3.
@@ 5.1 Krylov complexity
Figure 4: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
@@ 5.2 Krylov Entropy
Figure 5: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
@@ 6.1.1 $b_{n}$ and $\mu_{2}$
Figure 6: The plot of $b_{n}=\lambda$ with $n=2$.The numerical of $b_{n}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
@@
Figure 7: The numerical of dissipative coefficient $\mu_{2}$ in terms of $\log_{10}a$ for RD and MD, where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity. The value of $\mu_{2}$ indicates the RD and MD are weak dissipative systems.
@@ 6.2 General discussion of $\phi_{n}$
where we have defined $\tilde{c}_{n}=n\chi\mu$ and $b_{n}=n\alpha$. As $\eta$ approaches infinity, the solution of Eq. (52) has the form $\phi(n)\propto e^{-n\xi}$ with $\xi=\frac{2\alpha}{\mu\chi}$, where this form will appear in RD or MD due to their weak dissipative nature. In the weak dissipative region, the corresponding Krylov complexity behaves as follows: it increases exponentially at the onset, with the average position given by $C_{k}\propto e^{2\alpha\eta}$ before reaching $\eta_{*}$. After this point, the Krylov complexity saturates to a constant value $\xi$ when $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln{\frac{2\alpha}{\chi\mu}}$ dubbed as the scrambling time. It is evident that $\eta_{*}$ determines the range of exponential growth of Krylov complexity, determined by the dynamical system itself. It is straightforward to demonstrate that $\exp(2a\eta_{*})=\xi$. After some simple algebra, we could obtain the formula $\xi$ in our case as follows,
@@ 6.3 Wave function within the method of open system
$\sum_{n=0}^{\infty}M_{n}(x;\delta,\eta_{1})\frac{t^{n}}{n!}=((1+\delta t)^{2}+t^{2})^{-\eta/2}\exp(x\arctan(\frac{t}{1+\delta t}))$ (58)
@@
$|\mathcal{O}(t))=e^{i\mathcal{L}\eta}|e_{0})=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\sum_{n=0}^{\infty}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(-1\exp(2i\phi_{k}(\eta))\tanh\eta)^{n}}{(1+\mu_{2}\tanh\eta)^{n}}|e_{n})$
@@ 6.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open system’s method
Figure 8: The numerical of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
@@ 6.6 Krylov complexity and Krylov entropy with the approach of open system
Figure 9: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
Figure 10: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.